x^2dy/dx-3y^4=2xy


1
Expert's answer
2022-06-13T23:02:13-0400
"x^2\dfrac{dy}{dx}-3y^4=2xy"

Let "u=y^3." Then "u'=3y^2y'"

"3x^2y^2y'-9y^6=6xy^3"

"x^2u'-6xu=9u^2"

"u'-\dfrac{6}{x}u=\dfrac{9}{x^2}u^2"

Bernoulli Equation


"z=u^{1-2}=u^{-1}"

"z'=-u^{-2}u'"

"z'+\dfrac{6}{x}z=-\dfrac{9}{x^2}"

"x^6z'+6x^5z=-9x^4"

"d(x^6z)=-9x^4dx"


"x^6z=-\dfrac{9x^5}{5}+\dfrac{C}{5}"

"z=\dfrac{C-9x^5}{5x^6}"

"y^3=\dfrac{5x^6}{C-9x^5}"